OpenStudy (anonymous):

final exam 1st question: Answer

OpenStudy (anonymous):

I think for part a the vout will be 5k*iA so by calculating iA is enough, what say guys.? for Part b its 5k*iN2 and part c 5k*(iA+iN2) If you get please tell, i am correct or not.

OpenStudy (anonymous):

try to simulate the circuit using multisim.

OpenStudy (anonymous):

Assuming you hav R=6k and rm=6k q1)use the formulae iA=2*iN1*R/(2*R+rm) =2*1*6/([2*6]+6)=2/3 Vout= iA*rm = 4 q2)use the formulae iA =(-iN2)*R/(2*R+rm)=-1/3 Vout= iA*rm= -2 q3) Ans = q1+q2 =2 q4) R+R =12

OpenStudy (anonymous):

ok. thank u so much. n even need help for Q4 n Q2.if u hav done jus post the formulas plz.

OpenStudy (anonymous):

The elements in the circuit have the following values: R=5kΩ and rm=15kΩ i got one try left i messed up any help

OpenStudy (anonymous):

@hejå 6 -3 3 10

OpenStudy (benzo):

(-iN2)*R/(2*R+rm)*rm=0 .....what im doing wrong

OpenStudy (anonymous):

Its not an equation. Just put in the values fpr iN2,R,rm ...and i got iA=-1/3 ..the answer Vout = iA*rm

OpenStudy (anonymous):

Pls note..The questions are in the order q1)Assuming that iN1=1mA and iN2=0A, what is the value of vOUT in Volts? q2)Assuming that iN1=0A and iN2=1mA, what is the value of vOUT in Volts? q3)Assuming that iN1=1mA and iN2=1mA, what is the value of vOUT in Volts? q4)Assuming that Input 2 is left as an open circuit, what is the Thevenin equivalent resistance (in kOhms) seen from Input 1? Answers posted above.

OpenStudy (anonymous):

Good night and thanks openstudy !

OpenStudy (anonymous):

pls i need formular for q6. these are my values. Assume that in the circuit above, V0=10V, T=0.5ms, Rw=1Ω, Lw=0.5mH and RL=4Ω.